OpenStudy (anonymous):

Which of the following pairs of slopes are of perpendicular lines? 1/2 and 2 3/4 and -4/3 -2/5 and -5/2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$m=\frac{-1}{m'}$$ where m and m' are perpendicular

OpenStudy (anonymous):

3/4 and -4/3 is the answer, im not sure how, i just guessed but i got it right

OpenStudy (skullpatrol):

what is the reciprocal of 3/4?

OpenStudy (anonymous):

4/3

OpenStudy (skullpatrol):

Correct, now the formula above says take the opposite of that and that is the slope of the perpendicular.

OpenStudy (anonymous):

the opposite of 4/3 is -4/3?

OpenStudy (skullpatrol):

Yes :)

OpenStudy (skullpatrol):

So, $$\frac{3}{4}$$and $$\frac{-4}{3}$$ are slopes of lines that are perpendicular.

OpenStudy (anonymous):

thank you(:

OpenStudy (skullpatrol):

Thank you for trying :)